# Post-hoc Popularity Bias Correction in GNN-based Collaborative Filtering

### v1

@@ 6. Conclusion
In this work, we propose a popularity debiasing method in a post-hoc manner, a new method to mitigate popularity bias in GNN-based CF by estimating interaction-level popularity and leveraging it to derive a popularity direction to remove popularity-driven components from node embeddings. Unlike existing approaches, our method requires no retraining and can be applied directly to pre-trained models. Extensive experiments on three real-world datasets show that our method consistently outperforms state-of-the-art baselines, improving both head and tail recommendations. We further demonstrate that our method generalizes to other GNN-based CF backbones. Future research can extend post-hoc methods to other forms of bias correction, such as exposure bias or position bias, and explore applications beyond CF, including content-based, hybrid, and dynamic recommendation settings.
@@ A.1. PPD Algorithm
8:  Compute bias score $b_{ui}$ using Eq.  (7)
@@
Table 6. A comparison of different methods across datasets using LightGCN as the backbone for HR@20, evaluated on the top $\bm{20\%}$ and bottom $\bm{80\%}$ of items sorted by interaction frequency in the training data. The best-performing method is shown in bold, and the second-best method is underlined. Performance improvement (%) is given compared to the best baseline.
Debiasing type Method KuaiRec Coat Yahoo! R3
Top 20% Bottom 80% Top 20% Bottom 80% Top 20% Bottom 80%
Post-hoc PPD 0.9043 0.0037 0.4840 0.3945 0.3632 0.0493
DAP (Chen et al., 2024) 0.3007 0.0000 0.4799 0.3125 0.3283 0.0454
None LightGCN (He et al., 2020) 0.0709 0.0000 0.4812 0.1641 0.3277 0.0440
In-training APDA (Zhou et al., 2023) 0.1099 0.0000 0.4309 0.1836 0.2989 0.0475
NAVIP (Kim et al., 2022) 0.0064 0.0000 0.4521 0.1406 0.3418 0.0441
MACR (Wei et al., 2021) 0.1872 0.0000 0.2766 0.2109 0.0030 0.0246
IPSCN (Joachims et al., 2017) 0.0060 0.0015 0.4787 0.3828 0.3034 0.0447
SAM-REG (Boratto et al., 2021) 0.1113 0.0034 0.4468 0.3892 0.3134 0.0189
PPAC (Ning et al., 2024) 0.6177 0.0000 0.4834 0.2227 0.2178 0.0452
Improvement (%) +46.4% +8.8% +0.1% +1.4% +6.3% +3.8%
@@
[…]
A.2. Computational Complexity Analysis
The computational complexity of our proposed PPD method is analyzed as follows. The global preference estimation using equation (5) computes similarities between each item and all users, which requires $O(|\mathcal{U}||\mathcal{I}|d)$ operations, where $d$ is the embedding dimension. The personalized preference estimation using equation (6) averages similarities between interacted items for each user, costing $O\!\left(\sum_{u\in\mathcal{U}}|N_{u}|^{2}d\right)$, which on average becomes $O(|\mathcal{E}|\bar{D}d)$, where $\bar{D}$ is the average degree of nodes. The popularity score computation using equation (7) requires $O(|\mathcal{E}|)$. For each node, computing the popularity and preference centroids using equations (8)–(9) involves aggregating embeddings from its neighbors, costing $O(|\mathcal{E}|d)$. Constructing the popularity direction vector for all nodes with equation (10) incurs an additional $O(|\mathcal{V}|d)$. Finally, projecting each embedding onto its popularity direction using equations (4.2)–(4.2) requires two dot products and one scalar–vector multiplication per node, while removing the projected component requires a vector subtraction, resulting in a complexity of $O(|\mathcal{V}|d)$.
The total time complexity of PPD is $O\big(|\mathcal{U}||\mathcal{I}|d\;+\;|\mathcal{E}|\bar{D}d\;+\;|\mathcal{E}|\;+\;|\mathcal{V}|d\big).$ Since $|\mathcal{U}|+|\mathcal{I}|=|\mathcal{V}|$, we have $|\mathcal{V}|\ll|\mathcal{U}||\mathcal{I}|$, and therefore the term $|\mathcal{V}|d$ is negligible compared to $|\mathcal{U}||\mathcal{I}|d$. In practice, recommender graphs are typically sparse, i.e., $|\mathcal{E}|\ll|\mathcal{U}||\mathcal{I}|$. Thus, the $|\mathcal{E}|$ term is also negligible. The complexity can therefore be simplified to $O\big(|\mathcal{U}||\mathcal{I}|d\;+\;|\mathcal{E}|\bar{D}d\big).$ Moreover, in sparse graphs, we usually have $\bar{D}\ll|\mathcal{I}|$ and the typical average degree $\bar{D}$ can also be small due to the long-tailed distribution. As a result, the dominating term in practice is $O(|\mathcal{U}||\mathcal{I}|d).$ In practice, this cost can be further reduced by randomly sampling a subset of users during the similarity computation for global preference for each item, resulting in $O(|\mathcal{U}_{s}||\mathcal{I}|d),$ where $|\mathcal{U}_{s}|$ denotes the number of sampled users. The existing post-hoc popularity bias correction method (Chen et al., 2024) requires clustering similar nodes using the $K$-means algorithm and, within each cluster, estimating the means of high-degree and low-degree nodes for each node. Specifically, $K$-means clustering can be computationally expensive, especially for large values of $K$.
Appendix B Experimental Details
B.1. Baselines
We evaluate the performance of our method relative to popularity bias correction approaches used in CF. We provide a description for each of the baselines.
  • MACR: The Model-Agnostic Counterfactual Reasoning (Wei et al., 2021) method uses multi-task learning and counterfactual inference to isolate and remove the direct effect popularity.
  • SAM-REG: The Sampling and Regularization (Boratto et al., 2021) method balances samples across popular and less-popular items and regularizes against relevance–popularity correlation to reduce popularity bias.
B.2. Additional Experimental Results
More results for RQ2. For question RQ2, we evaluate the performance of the PPD method in both popular and niche item groups. We present the results using Recall@20 and NDCG@20 metrics in the main body of the paper in Figure 3. Here, Table 6 shows that PPD consistently achieves the highest HR@20 across both head (top 20%) and tail (bottom 80%) items on all datasets. Compared to the strongest baseline, PPD improves HR@20 by +46.4% (head) and +233.3% (tail) on KuaiRec, +1.8% (head) and +10.1% (tail) on Coat, and +2.7% (head) and +9.1% (tail) on Yahoo! R3, demonstrating its effectiveness in improving recommendations for both popular and niche items.
More results for RQ3. For RQ3, we evaluate performance by varying the number of GNN layers, with results for Recall@20 shown in the main body of the paper in Figure 2. Here, we present NDCG@20 results in Figure 5, which exhibit similar trends. On KuaiRec, LightGCN peaks at two layers and then declines, while PPD continues to improve with additional layers. On Coat, both models achieve their best performance at two layers. For Yahoo! R3, PPD gets optimal results at two layers, whereas LightGCN slightly benefits from three layers. Overall, across all datasets and depths, PPD consistently outperforms LightGCN for NDCG@20 as well, demonstrating its robustness and mitigation of popularity bias.

### v2

@@ 6. Conclusion
In this work, we propose a popularity debiasing method in a post-hoc manner, a novel method to mitigate popularity bias in GNN-based CF by estimating interaction-level popularity and leveraging it to derive a popularity direction to remove popularity-driven components from node embeddings. Unlike existing approaches, our method requires no retraining and can be applied directly to pre-trained models. Extensive unbiased evaluations on three real-world datasets show that our method consistently outperforms state-of-the-art methods, improving both head and tail recommendations. We further demonstrate that our method generalizes to other GNN-based CF backbones. Future research can extend post-hoc methods to other types of bias, such as exposure bias or position bias, and explore applications beyond GNN-based CF, including other latent-factor models, as well as content-based and dynamic recommendation settings.
Acknowledgments
This research was partly supported by the National Science Foundation under Award No. 2217023.
@@ A.1. PPD algorithm
8:  Compute popularity bias score $b_{ui}$ using Eq.  (7)
@@
Method KuaiRec Coat Yahoo! R3
Recall NDCG HR Recall NDCG HR Recall NDCG HR
PPD 0.1352 0.3612 0.9978 0.2437 0.2176 0.6121 0.1381 0.0601 0.2031
Improvement +1017.4% +1811.1% +287.6% +5.8% +7.3% +5.6% +8.1% +3.1% +6.4%
Method KuaiRec Coat Yahoo! R3
Recall NDCG HR Recall NDCG HR Recall NDCG HR
PPD 0.0988 0.2899 0.9043 0.2407 0.1957 0.5871 0.1505 0.0695 0.2237
Appendix B Experimental Details
B.1. Baselines
We evaluate our method against popularity bias correction baselines in CF. We provide a description for each of the baselines.
  • MACR: The Model-Agnostic Counterfactual Reasoning (Wei et al., 2021) uses multi-task learning and counterfactual inference to isolate and remove the direct effect popularity.
  • SAM-REG: The Sampling and Regularization (Boratto et al., 2021) method balances samples across popular and less popular items and regularizes against relevance–popularity correlation to mitigate popularity bias.
B.2. Implementation details
Since LightGCN (He et al., 2020) is the most widely used backbone for GNN-based CF, we adopt it for all baselines and our method, unless otherwise specified. The model is implemented in PyTorch and optimized using Adam. We apply early stopping with a patience of $100$ epochs, stopping if Recall@$20$ on the validation set does not improve for $100$ consecutive epochs. This setting allows the model to achieve its best performance while maintaining the early stopping criterion. We tune the learning rate among $1{e}^{-3},1{e}^{-4},1{e}^{-5}$ and the regularization coefficient among $0.0,1{e}^{-3},1{e}^{-5}$. Batch size is 2,0482{,}048 and embedding dimension is $256$. All other parameters follow the default settings of LightGCN. Unless stated otherwise, we use a two-layer GCN that integrates up to two-hop neighbors. Popularity scores ($b_{ui}$) are computed by min–max normalizing global preference ($p_{i}$) and personalized preference ($r_{ui}$). Hyperparameters for our method are tuned with popularity penalty coefficient $\beta\in\{0.0,0.1,0.2,0.3\}$ and preference centroid coefficient $\phi\in\{0.0,0.25,0.50,0.75,1.0\}$. $\beta$ is restricted up to 0.3 to avoid over-penalizing popularity, which could suppress preference signals and weaken the estimation of personalized preference. The small constant $\epsilon$ in equations (8) and (9) is set to $1{e}^{-8}$. For unbiased validation, we split the unbiased test data of each original dataset, using one-third for validation and the remaining two-thirds for testing for KuaiRec and Coat. However, for Yahoo! R3, the unbiased test set contains a very small number of interactions on average. Therefore, we do not split it and directly use it as the test set. The validation dataset for Yahoo! R3 is constructed by sampling interactions from the training set. We use the dot product as the ${sim}(\cdot)$ function in equations (5) and (6). To validate our method on other GNN-based CF models, we also use SGL (Wu et al., 2021) as a backbone (see RQ4). Following (Zhang et al., 2023a), in our experiments, we set the contrastive loss temperature to $0.15$ and its weight to $0.2$ for SGL, while keeping all other settings at the default LightGCN settings.
@@ B.2. Implementation details
(a) KuaiRec dataset
(b) Coat dataset
(c) Yahoo! R3 dataset
B.3. Additional experimental results
Performance of PPD with SGL as backbone (RQ4). To validate the generalizability of our method, we apply PPD to another widely used GNN-based CF backbone, Self-supervised Graph Learning (SGL) (Wu et al., 2021). SGL augments LightGCN with self-supervised contrastive learning, generating multiple graph views via node dropout, edge dropout, and random walks. By maximizing agreement between views of the same node, SGL enhances representation robustness and recommendation accuracy. Evaluating PPD on SGL tests whether our debiasing strategy benefits other GNN-based CF backbones. Table 4 shows results on KuaiRec, Coat, and Yahoo! R3. On KuaiRec, PPD achieves $1017.4\%$ Recall@20, $1811.1\%$ NDCG@20, and $287.6\%$ HR@20 over SGL. For Coat, improvements are $5.8\%$ Recall@20, $7.3\%$ NDCG@20, and $5.6\%$ HR@20, while on Yahoo! R3, gains are $8.1\%$ Recall@20, $3.1\%$ NDCG@20, and $6.4\%$ HR@20. These results highlight two key findings. First, PPD shows substantial improvements on datasets with strong popularity bias, such as KuaiRec, where SGL alone struggles to capture unbiased signals. Second, even on less biased datasets like Coat and Yahoo! R3, PPD still yields consistent improvements. This demonstrates that PPD is not limited to a specific backbone but can be effectively applied on top of stronger backbones like SGL, further boosting recommendation accuracy through mitigating popularity bias. SGL is considered a more effective method than LightGCN. PPD improves performance across both backbones, even with the weaker LightGCN embeddings. This indicates that our popularity direction remains stable even with weaker embeddings and improves performance in both weaker and stronger recommendation models.
Impact of incorporating personalized preference into popularity score estimation (RQ5). In RQ5, we examine the role of personalized preference ($r_{ui}$) in estimating the popularity score ($b_{ui}$). To this end, we introduce a variant, PPD w/o-PP, which computes the popularity score solely based on global preference ($p_{i}$) without accounting for user-specific preferences. The results in Table 5 show that PPD consistently outperforms PPD w/o-PP across all datasets and evaluation metrics. This comparison reveals that relying only on global preference can be insufficient, as popular items may still align with true user interests (Zhao et al., 2022). Without personalized preference, the popularity score ($b_{ui}$) risks suppressing relevant popular items, conflating popularity-driven interactions with true user preferences. This is also evident from the NAVIP (Kim et al., 2022) method, which relies only on global preferences (from total interactions), and PPD clearly outperforms it (see Table 2). These observations demonstrate that accounting for personalized preference in estimating the popularity score is important, as it enables PPD to more accurately estimate the true popularity effect, enabling better performance.
Effect of the popularity penalty coefficient ($\bm{\beta}$) on PPD (RQ7). In RQ7, we evaluate the effect of the popularity penalty coefficient ($\beta$), which controls the strength of penalizing using global preference when estimating personalized preferences. As shown in Figure 4, performance varies across datasets and values of $\beta$. On KuaiRec, Recall@20 improves as $\beta$ increases up to 0.2 but declines when raised to 0.3, indicating that moderate global preference penalization yields the best personalized preference and, consequently, the highest performance. On Coat, the highest performance occurs at $\beta=0.1$, with larger values leading to reduced Recall@20. In contrast, Yahoo! R3 achieves its peak performance at $\beta=0.3$. These results highlight two key insights. First, without any penalty ($\beta=0.0$), the model cannot achieve its best performance, demonstrating the necessity of incorporating a popularity penalty in estimating personalized preferences. Second, the optimal value of $\beta$ is dataset-dependent, indicating that it should be tuned for each dataset to achieve optimal performance.